OpenStudy (anonymous):

what is 50+What=5(10+6

OpenStudy (anonymous):

30, because 10+6= 16 16*5= 80 50+30=80

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

Your Welcome :) if you can like my answer as best response :)

OpenStudy (anonymous):

kay i just did, can u help with another? (42+14)=what(6+2)

OpenStudy (anonymous):

It's 7 because 42+14= 56 and 6+2=8 so you divided 56/8 which is 7

OpenStudy (anonymous):

thanks so much:)

OpenStudy (anonymous):

:)